I am trying to find the minimum for f(x)=5x^2-20x+3. i think to find the min. you do -b/2a. In which case i did (-)-20/2(5) and got 2. My teacher said that this answer was incorrect. If you could please tell me where i went wrong it would be much appreciated. thanks!
Answer by KnightOwlTutor(293)

You are trying to get the vertex. In order to do this you need to complete the square f(x)=5x^2-20x+3.Set f(x)=0
0=5x^2-20x+3
Subtract 3 from both sides
-3=5x^2-20x
Factor out the 5 on the right hand side of the equation
-3=5(x^2-4x
divide the 4 by 2 and then square it.
-3=5(x^2-4x+4)
since you added an additonal 20 on the right side -don't forget to multiply the 4 you just added by 5
-3+20=5(x-2)^2
17/5=(x-2)^2
take the square root of each side 17/5sq rt=x-2
sqrt85/5=x-2
add two to both sides
x=2sqrt85/5